# Modulational instability of Geodesic-Acoustic-Mode packets

### v1

@@ 1 Introduction
The Geodesic Acoustic Mode (GAM) is a plasma oscillation observed in fusion reactors with toroidal geometry, such as tokamaks or stellarators. It develops when the $E\!\times\!B$ drift velocity of the zonal flows (ZFs) varies so strongly along the poloidal coordinate that the corresponding $m=\pm 1,n=0$ flow divergence (where $m$ and $n$ are the poloidal and toroidal mode numbers, respectively) cannot be compensated by parallel flows [1]. As a result, a characteristic $m=\pm 1,n=0$ “up-down antisymmetric” pressure mode emerges, which leads to an oscillation of the ZFs, i. e. the GAM [1, 2, 3]. GAMs are thus recognized to be the non-stationary branch of the zonal flows [1, 4] and their associated electric potential is (to the leading order) an $m=n=0$ structure. The name GAM stems from the geodesic magnetic field line curvature, which is responsible for plasma compressibility and thus a necessary condition for the emergence of the characteristic pressure mode.
@@
It was recently shown in Ref. [9] that in the regime of moderate nonlinearity the dynamics of undamped isolated GAM packets is well described by a (cubic) nonlinear Schrödinger equation (NLSE). The NLSE is a standard model [10] for describing nonlinear dispersive oscillations with a (linear) dispersion relation of the form
@@
which in the limit $k_{r}^{2}\rho_{i}^{2}\ll 1$ (where $\rho_{i}$ denotes the ion Larmor radius and $k_{r}$ is the radial wavevector associated with the GAM radial electric field) approximates to the standard gyrokinetic GAM dispersion relation [11], as will be discussed in sec. 2.1. Adopting an NLSE as a model equation for GAM packets is furthermore supported by the general result that plasma eigenmodes (more exactly, their radial envelope), arising in toroidal systems as a consequence of various types of instabilities, can be described by a nonlinear Schrödinger equation with integro-differential coefficients [12, 13, 14, 15]. The NLSE model has been applied and studied extensively in the contexts of deep water waves, light traveling through optical fibres, Bose-Einstein condensates and others [16, 17, 18]. Some predictions of the NLSE are well-known, like the emergence of solitons, nonlinear wave breaking, the nonlinear phase shift and susceptibility to the modulational instability (MI). While some of these phenomena have already been observed in gyrokinetic simulations of GAMs in Ref. [9], thus confirming the NLSE as a valid description of the dynamics, in this report the focus is set on the MI, which to the best of the authors’ knowledge has not yet been studied in the context of GAMs.
@@ 2.2 Dynamics
Anomalous Dispersion ${\cal G}>0$ $\tau_{e}\lesssim 5.45,$
No Dispersion ${\cal G}=0$ $\tau_{e}\approx 5.45,$
Normal Dispersion ${\cal G}<0$ $\tau_{e}\gtrsim 5.45.$
@@ 3.1 Introduction and Behaviour
When the perturbation wavevector $k_{\text{pert}}$ lies in the unstable range, which will be specified in sec. 3.2, due to the previously mentioned self-focusing effect the sinusoidal perturbation will grow exponentially (as long as the perturbation amplitude $a_{1}$ is small compared to the plane wave amplitude $a_{0}$) at the expense of the plane wave ($k_{r}=0$) component of the wave, which in the following will be called the wave background. More details on the growth process are given in App. B. When the perturbation has grown so large that it dominates the shape of the wave $\psi$ it saturates and acquires a large nonlinear phase shift (as mentioned in sec. 2.2) compared to the background, leading to strongly incoherent phase fronts. In this saturation phase the radial spectrum of the wave contains many high-$k_{r}$ components. When the nonlinear phase shift is large enough that the perturbation has skipped an entire oscillation compared to the background, the wave front reconnects and the perturbation decreases again. Finally, the initial condition is restored and the MI process can start anew, leading to a cyclic behaviour called Akhmediev Breathers [32]. A single Akhmediev Breather cycle together with the corresponding radial spectrum is shown in fig. 4.
@@ 3.2 Conditions for instability
where $k_{\text{lim}}$ marks the boundary between stable and unstable perturbation wavevectors. The corresponding growth rate, which will be labelled $\gamma_{\text{MI}}:=|\mbox{Im}\,\omega_{\text{pert}}|$ in the following, reaches its maximum value at the wavevector $|k_{\text{max}}|:=a_{0}\sqrt{2\alpha_{\text{NL}}/{\cal G}}$. From eq. (16) one finds that
@@ 4.1 Split-step NLSE solver
where ${.}\mathfrak{F}\left\{.\right\}$ is the Fourier transform which is calculated using the Fast-Fourier-Transform algorithm, and $\varphi_{\text{NL}}^{t}\left[.\right]$ and $\varphi_{\text{D}}^{t}\left[.\right]$ denote the exact flows (i. e. the temporal evolution starting from an initial condition $\hat{\psi}_{0}(r)$) for the isolated nonlinear and isolated dispersive equation, respectively.
@@ 4.2 ORB5
The gyrokinetic theory reduces the 6D kinetic theory by one dimension by averaging over the gyromotion to obtain a 5D problem describing the dynamics of the guiding centre distribution function $f_{s}$ of each species $s$. The set of gyrokinetic equations describing the dynamics of $f_{s}$ can be constructed in different ways. The model implemented in ORB5 is derived from a gyrokinetic Lagrangian describing the particle motion in a magnetic field [39]. The time-symmetric Hamiltonian within the Lagrangian conserves energy automatically, leading to a model which is particularly useful for numerical simulations [23]. ORB5 provides numeric results that exhibit strong agreement with other gyrokinetic codes [24]. Its recent application to the dynamics of isolated GAMs is documented in Refs. [28, 29, 40, 41, 42, 43]. The interested reader is referred to these references for more details about the numerical model.
In order to create GAMs in the ORB5 simulations, an electric field $E_{r}$ is initialised in ORB5 via a perturbation $\delta n$ of the ion density $n_{0}$, which is related to the electric field as
$\delta n_{i}(r)\propto r\displaystyle{\partial\over\partial r}\left(rE_{r}(r)\right).$ (26)
@@ 5 Simulation results
GAM damping is known to decrease with rising safety factor $q_{s}$ [28]. In order to fulfill the assumption of weak damping (compared to the MI growth rate $\gamma_{\text{MI}}$, eq. (20)) that was posed in sec. 1, $q_{s}=15$ is chosen. In this regime the damping term derived by Qiu et al. in Ref. [27] is applicable as the assumptions of $1/q_{s}^{2}\ll k_{r}^{2}\rho_{i}^{2}\ll 1$ and $\frac{1}{2}\tau_{e}k_{r}^{2}\rho_{i}^{2}\ll 1$ are satisfied for all of the chosen perturbation wavevectors $k_{\text{pert}}$ (note that for $\rho_{i}/a_{\text{min}}=5.025\cdot 10^{-3}$ and $k_{r}=10\cdot 2\pi/a_{\text{min}}$ the value of $\frac{1}{2}\tau_{e}k_{r}^{2}\rho_{i}^{2}\approx 0.1$ and for $k_{r}=14\cdot 2\pi/a_{\text{min}}$, $\frac{1}{2}\tau_{e}k_{r}^{2}\rho_{i}^{2}\approx 0.3$). We remark that the damping rate derived by Sugama and Watanabe, see Ref. [25], is not applicable due to the approximation $k_{r}^{2}\rho_{i}^{2}\ll 1/q_{s}^{2}$ employed there, which is not fulfilled for any of the chosen perturbation wavevectors $k_{\text{pert}}$. One obtains from Ref. [27] that the unmodulated packet and wavevectors $k_{r}<5\,\frac{2\pi}{a_{\text{min}}}$ are undamped (see fig. 10). However, damping rates of larger wavevectors, e. g. $\gamma_{\text{Qiu}}(k_{r}=10\,\frac{2\pi}{a_{\text{min}}})=-6.07\cdot 10^{-6}\,\omega_{ci}$ are in similar orders of magnitude as the MI growth rate $\gamma_{\text{MI}}(k_{r}=10\,\frac{2\pi}{a_{\text{min}}})=3.11\cdot 10^{-5}\,\omega_{\text{ci}}$, i. e. $|\gamma_{\text{Qiu}}|\approx\frac{1}{5}|\gamma_{\text{MI}}|$, which leads to the conclusion that damping may slow the growth of MI down but it is not expected to suppress MI.
@@ 5.3 Role of the perturbation wavelength
The aforementioned scheme, illustrated in fig. 11, is applied to all wavevectors $k=4,5,\dots,14\,\,\frac{2\pi}{a_{\text{min}}}$. In fig. 12, the MI growth rates obtained in this way are compared to the theoretical results of sec. 3 (see fig. 5 and eq. (19)), where a damping term is included according to eq. (30). The NLSE results clearly overestimate the GK MI growth rate $\gamma$ for $k_{\text{pert}}>6\,\frac{2\pi}{a_{\text{min}}}$, however, after adjusting the strength of the damping amplitude by multiplying it with a factor of 2.5, the simulation results show good agreement with the theoretic predictions, most notably for wavevectors in the domain $6\,\frac{2\pi}{a_{\text{min}}}\leq k_{\text{pert}}\leq 11\,\frac{2\pi}{a_{\text{min}}}$. This assumption is justified by the benchmark in Ref. [30] (see fig. 4b in this reference), where in a comparison of the GAM damping $\gamma_{\text{Qiu}}$ to numeric simulations a similar difference was observed.
@@ 6 Summary and conclusions
While the NLSE model proved to give accurate predictions of the general GAM behaviour, the exact dynamics shows significant differences. An improvement of the model, putting it on a more firm theoretical ground, requires a derivation of the NLSE equation for GAMs from first principles, which should be addressed in the near future.
Predictive capability to other parameter regimes is limited and will only improve with further studies on the unknown variable $\alpha_{\text{NL}}$.
Altogether, the results indicate that the possibility and impact of an MI on the GAM dynamics will be small. This conclusion becomes more evident by the fact that in this study, electrons were treated adiabatically in simulations and in the damping term, while recent research suggests that a kinetic treatment of the electrons will increase damping, thus decreasing the likeliness of MI even further [45, 46]. On the other hand, the self-focusing behaviour associated with the MI formation process is omnipresent in the regime of anomalous dispersion (${\cal G}>0$, $\tau_{e}\lesssim 5.45$) and may be observable in other simulations similarly to the case presented in sec. 5.6. The main reason behind this is the size of the involved structures and unstable wavevectors $k_{r}$: while the self-focusing effect only requires that a local maximum is present in the packet envelope, which can be of any size (or even simply the packet itself), MI formation requires that the maximum is “on top of” a relatively constant packet, thus demanding much finer structures and higher wavevectors. This is not only a much more unlikely initial condition to spontaneously develop in a tokamak, but is also much more significantly affected by the damping process illustrated in secs. 5.2-5.4.

### v2

@@ 1 Introduction
The Geodesic Acoustic Mode (GAM) is a plasma oscillation observed in magnetic-confinement fusion reactors with toroidal geometry, such as tokamaks or stellarators [1]. It develops because of the compressibility of the $E\!\times\!B$ drift velocity of the zonal flows (ZFs), which leads to $(m\!\!=\!\!\pm 1,n\!=\!0)$ pressure sidebands (where $m$ and $n$ are the poloidal and toroidal wavenumbers, respectively). As GAMs and ZFs satisfy charge neutrality, the perpendicular divergence of the diamagnetic current density is compensated by that of the polarization current, resulting in the GAM oscillation [2, 3, 4]. GAMs are thus recognized to be the non-stationary branch of the zonal flows [2, 5] and their associated electric potential is (to the leading order) an $m=n=0$ structure. The name GAM stems from the geodesic magnetic field line curvature, which is responsible for plasma compressibility and thus a necessary condition for the emergence of the characteristic pressure mode.
@@
Prior studies have established that as a direct consequence of nonlinear gyrokinetic theory, the GAM dynamics is well described by an equation of Schrödinger type – i. e. an equation whose linear contribution is exactly the linear Schrödinger equation, while the nonlinear dynamics neccitates an integro-differential expression, as has been discussed in sec. 6 of Ref. [3]. Indeed, this is not surprising as it is a general result that plasma eigenmodes (more exactly, their radial envelope), arising in toroidal systems as a consequence of various types of instabilities, can be described by a nonlinear Schrödinger equation with integro-differential coefficients [10, 11, 12, 13]. For the reduced case of an isolated GAM (i. e. without interaction with other modes) a similar model equation is provided in Ref. [14], where it was shown that in the regime of moderate nonlinearity the dynamics of undamped isolated GAM packets is well described by a cubic nonlinear Schrödinger equation (NLSE). The NLSE is a standard model [15] for describing nonlinear dispersive oscillations with a (linear) dispersion relation of the form
@@
which in the limit $k_{r}^{2}\rho_{i}^{2}\ll 1$ (where $\rho_{i}$ denotes the ion Larmor radius and $k_{r}$ is the radial wavevector associated with the GAM radial electric field) approximates to the standard gyrokinetic GAM dispersion relation [16, 29, 17], as will be discussed in sec. 2.1. The NLSE model has been applied and studied extensively in the contexts of deep water waves, light traveling through optical fibres, Bose-Einstein condensates and others [18, 19, 20]. Some predictions of the NLSE are well-known, like the emergence of solitons, nonlinear wave breaking, the nonlinear phase shift and susceptibility to the modulational instability (MI). While some of these phenomena have already been observed in gyrokinetic simulations of GAMs in Ref. [14], thus confirming the NLSE as a valid description of the dynamics, in this report the focus is set on the MI, which to the best of the authors’ knowledge has not yet been studied in the context of GAMs.
@@ 2.2 Dynamics
Anomalous Dispersion ${\cal G}>0$ $\tau_{e}\lesssim 5.45,$
No Dispersion ${\cal G}=0$ $\tau_{e}\approx 5.45,$
Normal Dispersion ${\cal G}<0$ $\tau_{e}\gtrsim 5.45.$
@@ 3.1 Introduction and Behaviour
When the perturbation wavevector $k_{\text{pert}}$ lies in the unstable range, which will be specified in sec. 3.2, due to the previously mentioned self-focusing effect the sinusoidal perturbation will grow exponentially (as long as the perturbation amplitude $a_{1}$ is small compared to the plane wave amplitude $a_{0}$) at the expense of the plane wave ($k_{r}=0$) component of the wave, which in the following will be called the background wave. More details on the growth process are given in App. B. When the perturbation has grown so large that it dominates the shape of the wave $\psi$ it saturates and acquires a large nonlinear phase shift (as mentioned in sec. 2.2) compared to the background, leading to strongly incoherent phase fronts. In this saturation phase the radial spectrum of the wave contains many high-$k_{r}$ components. When the nonlinear phase shift is large enough that the perturbation has skipped an entire oscillation compared to the background, the wave front reconnects and the perturbation decreases again. Finally, the initial condition is restored and the MI process can start anew, leading to a cyclic behaviour called Akhmediev Breathers [34]. A single Akhmediev Breather cycle together with the corresponding radial spectrum is shown in fig. 4.
@@ 3.2 Conditions for instability
where $k_{\text{lim}}$ marks the boundary between stable and unstable perturbation wavevectors. The perturbation wavevector $k^{2}_{\text{pert}}$ provides the frequency mismatch of the sideband wave with respect to the background (pump) wave. The corresponding growth rate, which will be labelled $\gamma_{\text{MI}}:=|\mbox{Im}\,\omega_{\text{pert}}|$ in the following, reaches its maximum value at the wavevector $|k_{\text{max}}|:=a_{0}\sqrt{2\alpha_{\text{NL}}/{\cal G}}$. From eq. (16) one finds that
@@ 4.1 Split-step NLSE solver
where ${.}\mathfrak{F}\left\{.\right\}$ is the spatial Fourier transform which is calculated using the Fast-Fourier-Transform algorithm, ${.}\mathfrak{F}^{-1}\!\left\{.\right\}$ denotes the inverse Fourier transform and $\varphi_{\text{NL}}^{t}\left[.\right]$ and $\varphi_{\text{D}}^{t}\left[.\right]$ denote the exact flows (i. e. the temporal evolution starting from an initial condition $\hat{\psi}_{0}(r)$) for the isolated nonlinear and isolated dispersive equation, respectively.
@@ 4.2 ORB5
The gyrokinetic theory reduces the 6D kinetic theory by one dimension by averaging over the gyromotion to obtain a 5D problem describing the dynamics of the guiding centre distribution function $f_{s}$ of each species $s$. The set of gyrokinetic equations describing the dynamics of $f_{s}$ can be constructed in different ways. The model implemented in ORB5 is derived from a gyrokinetic Lagrangian describing the particle motion in a magnetic field [41]. The time-symmetric Hamiltonian within the Lagrangian conserves energy automatically, leading to a model which is particularly useful for numerical simulations [25]. ORB5 provides numerical results that exhibit strong agreement with other gyrokinetic codes [26]. Its recent application to the dynamics of isolated GAMs is documented in Refs. [30, 31, 42, 43, 44, 45]. The interested reader is referred to these references for more details about the numerical model.
In order to create GAMs in the ORB5 simulations, an electric field $E_{r}$ is initialised in ORB5 via a perturbation $\delta n_{i}$ of the gyrocentre density of the only ion species included in the simulations. This perturbation leads through the gyrokinetic Poisson equation to a potential perturbation (or equivalently to a polarization density) that ensures quasineutrality. The relevant relation between density perturbation and electric field for a (0,0) perturbation can be expressed as
$\delta n_{i}(r)\propto{1\over r}\displaystyle{\partial\over\partial r}\left(rE_{r}(r)\right).$ (26)
@@ 5 Simulation results
GAM damping is known to decrease with rising safety factor $q_{s}$ [30]. In order to fulfill the assumption of weak damping (compared to the MI growth rate $\gamma_{\text{MI}}$, eq. (20)) that was posed in sec. 1, the safety factor $q_{s}=15$ is chosen. In this regime the damping term derived by Qiu et al. in Ref. [29] is applicable as the assumptions of $1/q_{s}^{2}\ll k_{r}^{2}\rho_{i}^{2}\ll 1$ and $\frac{1}{2}\tau_{e}k_{r}^{2}\rho_{i}^{2}\ll 1$ are satisfied for all of the chosen perturbation wavevectors $k_{\text{pert}}$ (note that for $\rho_{i}/a_{\text{min}}=5.025\cdot 10^{-3}$ and $k_{r}=10\cdot 2\pi/a_{\text{min}}$ the value of $\frac{1}{2}\tau_{e}k_{r}^{2}\rho_{i}^{2}\approx 0.1$ and for $k_{r}=14\cdot 2\pi/a_{\text{min}}$, $\frac{1}{2}\tau_{e}k_{r}^{2}\rho_{i}^{2}\approx 0.3$). We remark that GAM damping has also been studied under different approximation limits, e. g. by Sugama and Watanabe [27] whose expression is appropriate in the limit $k_{r}^{2}\rho_{i}^{2}\ll 1/q_{s}^{2}$, and by Gao et al. [46], who employ the approximation $\tau_{e}\rightarrow 0$. For a detailed discussion about the damping behaviour in the transitory regime between the limits by Refs. [29, 27, 46] the reader is referred to Ref. [16]. One obtains from Ref. [29] that the unmodulated packet and wavevectors $k_{r}<5\,\frac{2\pi}{a_{\text{min}}}$ are undamped (see fig. 10). However, damping rates of larger wavevectors, e. g. $\gamma_{\text{Qiu}}(k_{r}=10\,\frac{2\pi}{a_{\text{min}}})=-6.07\cdot 10^{-6}\,\omega_{ci}$ are in similar orders of magnitude as the MI growth rate $\gamma_{\text{MI}}(k_{r}=10\,\frac{2\pi}{a_{\text{min}}})=3.11\cdot 10^{-5}\,\omega_{\text{ci}}$, i. e. $|\gamma_{\text{Qiu}}|\approx\frac{1}{5}|\gamma_{\text{MI}}|$, which leads to the conclusion that damping may slow the growth of MI down but it is not expected to suppress MI.
@@ 5.3 Role of the perturbation wavelength
The aforementioned scheme, illustrated in fig. 11, is applied to all simulations, with respective wavevectors $k_{\text{pert}}=4,5,\dots,14\,\,\frac{2\pi}{a_{\text{min}}}$. In fig. 12, the MI growth rates obtained from the GK simulation data employing the method described above are compared to the theoretical results of sec. 3 (see fig. 5 and eq. (19)). The theoretical MI predictions, which do not contain damping, clearly overestimate the GK MI growth rate $\gamma$ for high perturbation wavevectors $k_{\text{pert}}>6\,\frac{2\pi}{a_{\text{min}}}$. As can be seen in fig. 12(a), even when the nominal damping rate following eq. (30) is included the growthrate of the MI is overpredicted as compared to the GK simulations. However, after adjusting the strength of the damping amplitude by multiplying it with a factor of 2.5, the simulation results show good agreement with the theoretic predictions (fig. 12(b)). Not only does this adjustment improve the matching of the threshold between stable and unstable perturbations at $k_{\text{pert}}\approx 11\,\frac{2\pi}{a_{\text{min}}}$, but it furthermore fits better to the values of the growthrates for the wavevectors in the domain $6\,\frac{2\pi}{a_{\text{min}}}\leq k_{\text{pert}}\leq 11\,\frac{2\pi}{a_{\text{min}}}$. We remark that this correction factor is also in line with the benchmark in Ref. [32] (see fig. 4b in this reference), where a similar difference between numeric simulations and theoretical prediction has been observed.
@@ 5.6 Self-focusing of Gaussian packets
One can observe that in all three simulations the Gaussian packet experiences self-focusing, resulting in an increase of the maximal amplitude and a phase skip of the Gaussian center compared to the packet edges. It is apparent that this behaviour is very similar to the MI, while it in contrast does not require an initial modulation of the envelope. We should note at this point that, differently from the in sec. 3.1 presented standard scenario of the MI, the background wave is not a plane wave but a packet with a finite spectrum. Hence, it contains also higher spectral components, which can act as a seed for MI. Indeed, a spectral analysis, which for the sake of brevity is not reported here, reveals exponential growth of the wavevectors $k_{r}=3\,\frac{2\pi}{a_{\text{min}}}$, $4\,\frac{2\pi}{a_{\text{min}}}$ and $5\,\frac{2\pi}{a_{\text{min}}}$. This observation suggests an interpretation of this behaviour as an MI, whereas the observed unstable wavevectors are not necessarily the ones with the strongest growth rates but also depend on the initial conditions of the packet. We should add that the generation of higher spectral components is an inherent part of the nonlinear dynamics of the GAM, which might contribute to the observed spectral behaviour on top of the MI.
@@ 6 Summary and conclusions
While the NLSE model proved to give accurate predictions of the general GAM behaviour, the exact dynamics shows significant differences. An improvement of the model, putting it on a more firm theoretical ground, requires a derivation of the NLSE equation for GAMs from first principles, which should be addressed in the near future. Predictive capability to other parameter regimes is limited and will only improve with further studies on the unknown variable $\alpha_{\text{NL}}$.
Altogether, the results indicate that (without considering interactions with other plasma modes) the possibility and impact of an MI on the GAM dynamics will be small. This conclusion becomes more evident by the fact that in this study, electrons were treated adiabatically in simulations and in the damping term, while recent research suggests that a kinetic treatment of the electrons will increase damping, thus decreasing the likeliness of MI even further [48, 49]. On the other hand, the self-focusing behaviour associated with the MI formation process is omnipresent in the regime of anomalous dispersion (${\cal G}>0$, $\tau_{e}\lesssim 5.45$) and may be observable in other simulations similarly to the case presented in sec. 5.6. The main reason behind this is the size of the involved structures and unstable wavevectors $k_{r}$: while the self-focusing effect only requires that a local maximum is present in the packet envelope, which can be of any size (or even simply the packet itself), MI formation requires that the maximum is “on top of” a relatively constant packet, thus demanding much finer structures and higher wavevectors. This is not only a much more unlikely initial condition to spontaneously develop in a tokamak, but is also much more significantly affected by the damping process illustrated in secs. 5.2-5.4. However, this conclusion is only valid in the absence of interaction with other plasma modes.